On January 1, 2017, a company issued 10-year, 10% bonds payable with a par value of $500,000 and received $442,647 in cash proceeds. The market rate of interest at the date of issuance was 12%. The bonds pay interest semiannually on July 1 and January 1. The issuer uses the straight-line method for amortization. Prepare the issuer's general journal entry to record the first semiannual interest payment on July 1, 2017.

a) Prepare the journal entry to issue the bond on January 1, 2017

Date

                    Explanation

Debit

Credit

  1. Prepare the journal entries to record the first two semiannual interest payments on June 30, 2017 and December 31, 2017

Date

                    Explanation

Debit

Credit

  1. Calculate the total bond interest expense for the year ended December 31, 2017.

                                   

  1. Calculate the bond carrying value as of December 31, 2017
Answer #1

a) Prepare the journal entry to issue the bond on January 1, 2017

Date

                    Explanation

Debit

Credit

Jan 1 Cash 442647
Discount on bonds payable 57353
Bonds payable 500000
  1. Prepare the journal entries to record the first two semiannual interest payments on June 30, 2017 and December 31, 2017

Date

                    Explanation

Debit

Credit

June 30 Interest expense 27867.65
Discount on bonds payable (57353/20) 2867.65
Interest payable (500000*5%) 25000
Dec 31 Interest expense 27867.65
Discount on bonds payable 2867.65
Interest payable 25000

c) Total bond interest expense = 27867.65*2 = 55735.30

d) Bond carrying value = 442647+2867.65+2867.65 = 448382.30